OpenStudy (anonymous):

Why does: y+dy=(x+dx) squared equal: y+dy=x squared +2x * dx +(dx) squared?

OpenStudy (anonymous):

You're just squaring a binomial: \[(a+b)^2=(a+b)(a+b)=a^2+2ab+b^2\]